Nick and Liz have decided to move from the city to the suburbs this means that they will have to make the essential purchase of a car in order to get to work they researched used 2 year old cars of the same make model and condition and equipped with the same options they found a website stating that the average price should be 18500 these are the prices they were quoted 15500,18800,16900,19900,18000,21000

if they continued their search for one more price quote what would that price have to be so that the mean of all seven of the car prices would be the same as the mean quotes on the website


1
Expert's answer
2021-08-22T17:04:38-0400

Mean= sum of terms/number of terms

Let the 7th term be x

therefore,

18500= (15500+18800+16900+19900+18000+21000+x)/7

18500 = (110100+x)/7

multiplying both sides by 7

18500*7=110100+x

129500=110100+x

x=129500-110100

x= 19400

The 7th price quote will be 19400